Investor has written ten call options for one car each with a strike price of 80,000 PLN per car and premium of 1% strike price. what would be investors result if the actual market price at delivery date is 90,000 PLNi
Answer #1

Strike price = PLN 80,000

Premium = 80,000 * 1% = 800

Actual market price at delivery date = PLN 90,000

Since, the price at the delivery date is higher, the investor will excercise the call option and his profit would be

Profit = 90,000 - 80,000 - 800 = 9,200 per car

The investor will get the benefit of 9,200 per car by excersing the call option.